Find two vectors of norm 1 that are orthogonal to the vectors u = (2,1, −4,0), v = (−1, −1,2,2),

and w = (3,2,5,4).


1
Expert's answer
2021-08-05T15:38:29-0400

Let vector "x=(a, b, c, d)" is orthogonal to the vectors "u=(2,1,-4,0), v=(-1,-1,2,2),"

and "w=(3,2,5,4)." Then


"\begin{matrix}
 2a+b-4c=0 \\
 -a-b+2c+2d=0 \\
 3a+2b+5c+4d=0 \\

\end{matrix}"

Augmented matrix


"A=\begin{pmatrix}
2 & 1 & -4 & 0 & & 0 \\
-1 & -1 & 2 & 2 & & 0 \\
3 & 2 & 5 & 4 & & 0 \\
\end{pmatrix}"

"R_1=R_1\/2"


"\begin{pmatrix}
1 & 1/2 & -2 & 0 & & 0 \\
-1 & -1 & 2 & 2 & & 0 \\
3 & 2 & 5 & 4 & & 0 \\
\end{pmatrix}"

"R_2=R_2+R_1"


"\begin{pmatrix}
1 & 1/2 & -2 & 0 & & 0 \\
0 & -1/2 & 0 & 2 & & 0 \\
3 & 2 & 5 & 4 & & 0 \\
\end{pmatrix}"

"R_3=R_3-3R_1"


"\begin{pmatrix}
1 & 1/2 & -2 & 0 & & 0 \\
0 & -1/2 & 0 & 2 & & 0 \\
0 & 1/2 & 11 & 4 & & 0 \\
\end{pmatrix}"

"R_2=-2R_2"


"\begin{pmatrix}
1 & 1/2 & -2 & 0 & & 0 \\
0 & 1 & 0 & -4& & 0 \\
0 & 1/2 & 11 & 4 & & 0 \\
\end{pmatrix}"

"R_1=R_1-R_2\/2"


"\begin{pmatrix}
1 & 0 & -2 & 2 & & 0 \\
0 & 1 & 0 & -4& & 0 \\
0 & 1/2 & 11 & 4 & & 0 \\
\end{pmatrix}"

"R_3=R_3-R_2\/2"


"\begin{pmatrix}
1 & 0 & -2 & 2 & & 0 \\
0 & 1 & 0 & -4& & 0 \\
0 & 0 & 11 & 6 & & 0 \\
\end{pmatrix}"

"R_3=R_3\/11"


"\begin{pmatrix}
1 & 0 & -2 & 2 & & 0 \\
0 & 1 & 0 & -4& & 0 \\
0 & 0 & 1 & 6/11 & & 0 \\
\end{pmatrix}"

"R_1=R_1+2R_3"


"\begin{pmatrix}
1 & 0 & 0 & 34/11 & & 0 \\
0 & 1 & 0 & -4& & 0 \\
0 & 0 & 1 & 6/11 & & 0 \\
\end{pmatrix}"

"x=(-\dfrac{34}{11}t, 4t, -\dfrac{6}{11}t, t), t\in \R"

"|x|=\dfrac{|t|}{11}\sqrt{(-34)^2+(44)^2+(-6)^2+(11)^2}=\dfrac{57}{11}|t|"


"t=\frac{11}{57}:\,x_1=(-\dfrac{34}{57}, \dfrac{44}{57}, -\dfrac{6}{57}, \dfrac{11}{57})"

"t=-\frac{11}{57}:\,x_2=(\dfrac{34}{57}, -\dfrac{44}{57}, \dfrac{6}{57}, -\dfrac{11}{57})"